What is the standard enthalpy of formation of     What is the standard enthalpy of formation of    

CH 3 CH 2 CH 2 CHO(l)? CH3CH2CH2CHO(l)? 2CH 3 CH 2 CH 2 CHO(l)+5O 2 (g)→8H 2 O(l)+8CO 2 (g); 2CH3CH2CH2CHO(l)+5O2(g)→8H2O(l)+8CO2(g);

ΔH°=–4943.6 kJ ΔH°=–4943.6 kJ

Substance ΔH° f (kJ/mol)

CO 2 (g) -393.5

H 2 O(l) –285.8   

a.
–245.4 kJ/mol

b.
+245.4 kJ/mol

c.
–1792.5 kJ/mol

d.
–3151.1 kJ/mol

e.
+3151.1 kJ/mol
Answer #1

delta(H) of reaction = delta(H f)product - delta(H f)reactant

delta(H f) of O2(g) = 0

-4943.6 = 8*(-393.5) + 8*(-285.8) - 5*0 - 2*x

-4943.6 = -3148 - 2286.4 -2*x

2*x = 4943.6  -3148 - 2286.4

2*x = 4943.6 - 5434.4

2*x = -490.8

x = -245.4 KJ/mol

standard enthalpy of formation of CH 3 CH 2 CH 2 CHO = -245.4 KJ/mol

Answer : option a